# Discrimination performance in illness-death models with interval-censored disease data

### v1

@@ 1 Introduction
The aim of this article is to study the discrimination performance of an interval-censored binary disease marker on survival. How much does the occurrence or absence of disease contribute to survival predictions over time? The illness-death model for data in which the disease state is interval-censored is considered. The effect of interval-censoring on the time-specific Area Under the receiver operating characteristic Curve (AUC) in both incident/dynamic and cumulative/dynamic definition is evaluated (Heagerty and Zheng, , 2005; Zheng and Heagerty, , 2007). Several estimation approaches are compared for two types of models: the Cox model with time-dependent disease marker and the illness-death model for interval-censored data as implemented in the msm and SmoothHazard R-packages (Jackson, , 2011; Touraine et al., , 2017). For this purpose a simulation study is conducted where data are simulated from an illness-death model with Weibull transition hazards and disease state censored at regular intervals. The different methods are also applied to soft tissue sarcoma data, where patients are initially disease-free after surgery, and the interval-censored disease state is the post-operative development of distant metastases.
@@ 2 Illness-death models
  • $V(t)=2$ corresponds to $T^{D}=t$ (for simplicity of notation, it is assumed that the subject has just died, i.e., $T^{D}>t-$), with two possible scenarios: (i) $X(t-)=0$ if the subject die directly, or (ii) $X(t-)=1$ if the subject first develop the disease and then die.
Often, the exact time of disease onset cannot be observed; it can only be diagnosed at pre-specified follow-up times, resulting in interval-censored data. In this article, four different methods to estimate the illness-death model data are compared: (1) the Cox, (1972) model with disease state as time-dependent covariate (ignoring the interval-censored nature of the time-dependent covariate), (2) the piecewise-constant illness-death model accounting for interval-censoring using the msm function from the msm package (Jackson, , 2024, 2011), (3) the Weibull illness-death model accounting for interval-censoring using the idm(method = "Weib") function from the SmoothHazard package, and (4) the M-spline illness-death model accounting for interval-censoring using the idm(method = "Splines") function from the SmoothHazard package (Touraine et al., , 2024, 2017). All methods are now introduced and then compared in terms of discrimination performance in a simulation study (Section 4) and in a real data application (Section 5).
@@ 2.1 Cox model with time-dependent covariate
where $\lambda_{0}(t)$ is the baseline hazard, $X(t)$ is the binary disease marker at time $t$ and $\beta$ its effect. This model can be estimated by e.g. the coxph R-function from the survival package (Therneau and Grambsch, , 2000), which ignores the interval-censored nature of the time-dependent covariate. Disease time is assumed to be the time of diagnosis of disease: $X(t)=0$ if a patient was not diagnosed with disease yet at time $t$ and $X(t)=1$ if a patient was diagnosed with disease by time $t$.
@@ 2.2 Piecewise-constant model accounting for interval-censoring
where $c_{k}$ are the times at which the hazard may change. This model is implemented in the msm package by Jackson, (2011) and can account for interval-censoring and the probability of developing disease between last disease scan and death or lost to follow-up.
The hazards for transitions to the death state can also be assumed proportional, allowing the effect of disease on survival to be modeled by constraining the transition hazards $\lambda_{12}$ and $\lambda_{02}$ to be equal. This can be done using the qconstrained option and by defining a binary time-dependent covariate $X(t)$, which equals 0 if the patient has not yet been diagnosed with the disease at time $t$, and 1 if the diagnosis has occurred by that time. The transition hazard from state 1 (disease) to state 2 (death) is then equal to $\lambda_{12}(t)=\lambda_{02}\exp\{\beta X(t)\}$.
@@ 3 Discrimination for time-dependent marker and survival data
To extend the concepts of sensitivity and specificity to settings with censored data and longitudinal markers, several definitions of cases and controls have been proposed. In this article, two of these definitions are considered: 1) incident cases and dynamic controls, and 2) cumulative cases and dynamic controls (Heagerty et al., , 2000; Heagerty and Zheng, , 2005; Zheng and Heagerty, , 2007).
@@ 3.1 Incident cases and dynamic controls
$p(t)=$ $\Pr\left(X_{i}(t-)=1\mid T^{D}_{i}=t\right)=\frac{\frac{\lambda_{12}(t)}{\lambda_{02}(t)}P_{01}(t-)}{P_{00}(t-)+\frac{\lambda_{12}(t)}{\lambda_{02}(t)}P_{01}(t-)},$ (11)
@@
$\hat{\pi}_{1}(t)=\frac{\hat{P}_{01}(t)}{\hat{P}_{00}(t)+\hat{P}_{01}(t)}\quad\text{ and }\quad\hat{p}(t)=\frac{\frac{\hat{\lambda}_{12}(t)}{\hat{\lambda}_{02}(t)}\hat{P}_{01}(t-)}{\hat{P}_{00}(t-)+\frac{\hat{\lambda}_{12}(t)}{\hat{\lambda}_{02}(t)}\hat{P}_{01}(t-)},$
@@ 3.2 Cumulative cases and dynamic controls
where $T^{D}$ is time of death, $X(s)$ is marker measurement at time $s$. Cases are individuals who die within a time window ($t-s$) from $s$ and controls are individuals who survive the time window. The cumulative/dynamic AUC is then defined by
@@
$\hat{\pi}_{1}(s,t)=\frac{\hat{P}_{11}(s,t)\hat{P}_{01}(0,s)}{\hat{P}_{00}(0,t)+\hat{P}_{01}(0,t)}\quad\text{ and }\quad\hat{p}(s,t)=\frac{\hat{P}_{12}(s,t)\hat{P}_{01}(0,s)}{\hat{P}_{02}(s,t)\hat{P}_{00}(0,s)+\hat{P}_{12}(s,t)\hat{P}_{01}(0,s)},$
@@ 4 Simulation study
To study the discrimination performance of an interval-censored binary disease marker on survival a simulation study was conducted. Incident/dynamic and cumulative/dynamic AUC were computed for the different estimation procedures presented in Section 2: the Cox model with time-dependent disease marker, which ignores interval-censoring, and the illness-death model with piecewise-constant, Weibull or M-spline transition hazards. For the piecewise-constant model, the four change points $c_{k}$ at which the hazard may change were considered at 6, 30, 60, and 90 months. For the M-spline model, the default of 7 knots per transition was used. Multiple data scenarios were simulated (Section 4.1) and results from the different methods were compared (Section 4.2).
@@ 4.1 Data generation and methods
Table 1 summarizes the twelve simulated scenarios. The number of individuals per data set, $N$, was either 1000 or 2000. The survival time was censored according to two different censoring schemes: either it was censored administratively at 10 years follow-up or censoring times were sampled from a uniform distribution between 5 and 10 years. The disease state was observed only at pre-specified follow-up visits. The scenarios cover three different follow-up schemes in which the disease state was observed every 3, 6, or 12 months. Data were generated from Weibull transition hazards with a common shape parameter $k=0.5$ and different rate parameters $\alpha_{01}=0.05,\alpha_{02}=0.05$ and $\alpha_{12}=0.56$. These parameters were kept fixed throughout the simulated scenarios and were based on the data discussed in Section 5, assuming a proportional hazards between diseased and non diseased with HR equal to 11.2. The true AUC values over time were determined by transition probabilities, derived from the Weibull transition hazards, as shown in Supplementary Material B.
@@
Each scenario was simulated $n_{sim}=1000$ times. For each model and scenario, estimated incident/dynamic and cumulative/dynamic AUCs over the years were presented graphically. Estimated bias, empirical standard error, and root mean squared error (Morris et al., , 2019) of both AUCs at specific time-points were considered as performance measures.
@@
Table 1: Simulated scenarios. $N$ is the total number of patients. Censoring is the type of death censoring: Unif(5, 10) means censoring was uniformly sampled between 5 and 10 years and 10 means that administrative censoring occurred at 10 years. Follow-up is time between disease observations in months.
Scenario $N$ Censoring Follow-up
@@ 4.2 Results
Hazard ratios for disease (yes vs no) were estimated for the piecewise-constant and the Cox model. Average values over $n_{sim}=1000$ repetitions were summarized in Supplementary Table C.1. For the Weibull and M-spline model no effect could be estimated, since the idm function does not allow transition hazards to be proportional. The coefficients from the Cox model were consistently more biased than the ones from the piecewise-constant model. The Cox model underestimated the true coefficient and the bias increased for larger follow-up intervals. These results are in line with Leffondré et al., (2013) who showed that the effect estimates of the Cox model were biased if the covariate affected both the risk of disease and death.
The M-spline model did not converge for many simulated data sets. The number of invalid estimations is shown in Supplementary Table C.2. Earlier censoring according to the uniform distribution between 5 and 10 years (scenarios A-C, G-I) resulted in a larger percentage of invalid estimations, compared to administrative censoring at 10 years (scenarios D-F, J-L). These invalid estimates prevented the estimation of the incident/dynamic and cumulative/dynamic AUCs. Therefore, the results for the M-spline model shown in Sections 4.2.2 and 4.2.3 are only based on the obtained valid estimates.
@@ 5.1 Soft tissue sarcoma data
Figure 4: Soft tissue sarcoma illness-death model ($N$ = 2232). State 0 (disease-free): surgery with curative intent; state 1 (disease): development of distant metastases (DM); state 2: death. The numbers indicate the number of patients moving from one state to the other.
@@ 5.2 Results
The HRs estimated by the Cox model and the piecewise-constant model were equal to 11.71 (95% CI = [10.31; 13.29]) and 11.28 (95% CI = [9.82; 12.96]) respectively. Figure 5 displays on the left panel the non-parametric cumulative baseline hazards and on the right panel a graphical check of their fit to a Weibull distribution. For this figure the time of DM was assumed to be equal to the time that DM was detected during screening. If the data came from a Weibull distribution, the plot of the log-cumulative hazard against the log of time should be approximately linear (right panel), which is not the case in particular for the transition from surgery to DM (black line). The Weibull model therefore may not be appropriate for this data. The M-spline model did not converge for this data set and consequently the discrimination performance could not be estimated.
@@
Figure 6 shows the AUCs over time for the different models (left panel: incident/dynamic; right panel: cumulative/dynamic), with time-specific values at years $t=1,2,3,4,5$ shown in Table 4. The incident/dynamic AUC of the Weibull model (orange) is initially much larger compared to the other models and declines over time. The incident/dynamic AUC of the piecewise-constant model (magenta) is the lowest of all three methods. The cumulative/dynamic AUC of the Cox model (blue) is generally the largest and the Weibull models the lowest.
@@
Cox ROC (green): estimate based on Cox model (4) through risksetAUC function (only for incident/dynamic AUC). Cox Prob (blue): estimate based on Cox model through mstate transition probabilities. PW-const (magenta): estimate based on piecewise-constant model (5) through msm. Weibull (orange): estimate based on Weibull model (6) through SmoothHazard.
@@
Time-specific AUC Year $t$ Cox ROC Cox Prob PW-const Weibull
$\widehat{\text{AUC}}^{\text{I/D}}(t)$ 1 0.74 0.75 0.71 0.81
2 0.76 0.76 0.73 0.78
3 0.76 0.76 0.73 0.76
4 0.75 0.75 0.71 0.74
5 0.74 0.74 0.70 0.72
$\widehat{\text{AUC}}^{\text{C/D}}(t,t+5)$ 1 0.64 0.62 0.62
2 0.69 0.66 0.63
3 0.70 0.68 0.64
4 0.68 0.66 0.64
5 0.67 0.66 0.64
Note: Cox ROC, estimate based on Cox model (4) through risksetAUC function (only for incident/dynamic AUC); Cox Prob, estimate based on Cox model through mstate transition probabilities; PW-const, estimate based on piecewise-constant model (5) through msm; Weibull, estimate based on Weibull model (6) through SmoothHazard.

### v2

@@ Abstract
Disclaimer: This is the accepted version of the manuscript that has been published in Statistical Methods in Medical Research. Please cite the published article available at DOI: 10.1177/09622802251412855.
@@ 1 Introduction
The aim of this article is to study the discrimination performance of an interval-censored binary disease marker on survival. How much does the occurrence or absence of disease contribute to survival predictions over time? The illness-death model for data in which the disease state is interval-censored is considered. The effect of interval-censoring on the time-specific Area Under the receiver operating characteristic Curve (AUC), for both incident/dynamic and cumulative/dynamic definition (Heagerty and Zheng, , 2005; Zheng and Heagerty, , 2007), is addressed by extending the standard time-dependent AUC to incorporate hazards and transition probabilities. This allows the AUC calculation to account for the uncertainty inherent in interval-censored data, which is ignored by traditional methods assuming exact times. Several estimation approaches are compared for two types of models: the Cox model with time-dependent disease marker and the illness-death model for interval-censored data as implemented in the msm and SmoothHazard R-packages (Jackson, , 2011; Touraine et al., , 2017). For this purpose a simulation study is conducted where data are simulated from an illness-death model with Weibull transition hazards and disease state censored at regular intervals. The different methods are also applied to soft tissue sarcoma data, where patients are initially disease-free after surgery, and the interval-censored disease state is the post-operative development of distant metastases.
@@ 2 Illness-death models
  • $V(t)=2$ corresponds to $T^{D}=t$ (for simplicity of notation, it is assumed that the subject has just died, i.e., $T^{D}>t$-), with two possible scenarios: (i) X(t)=0X(t-)=0 if the subject die directly, or (ii) X(t)=1X(t-)=1 if the subject first develop the disease and then die.
Often, the exact time of disease onset cannot be observed; it can only be diagnosed at pre-specified follow-up times, resulting in interval-censored data. In this article, four different methods to estimate the illness-death model data are compared: (1) the Cox, (1972) model with disease state as time-dependent covariate (ignoring the interval-censored nature of the time-dependent covariate), (2) the piecewise-constant illness-death model accounting for interval-censoring using the msm function from the msm package (Jackson, , 2024, 2011), (3) the Weibull illness-death model accounting for interval-censoring using the idm(method = "Weib") function from the SmoothHazard package, and (4) the M-spline illness-death model accounting for interval-censoring using the idm(method = "Splines") function from the SmoothHazard package (Touraine et al., , 2024, 2017). All methods are now introduced and then compared in terms of discrimination performance in a simulation study (Section 4) and in a real data application (Section 5). Each model is presented without covariate adjustment, as covariates are not included in the simulation study or in the application, but they could be incorporated on a transition-specific basis to capture distinct effects.
@@ 2.1 Cox model with time-dependent covariate
where $\lambda_{0}(t)$ is the baseline hazard, $X(t)$ is the binary disease marker at time $t$ and $\beta$ its effect. This model can be estimated by e.g. the coxph R-function from the survival package (Therneau and Grambsch, , 2000), which ignores the interval-censored nature of the time-dependent covariate. Disease time is assumed to be the time of diagnosis of disease: $X(t)=0$ if a patient was not diagnosed with disease yet at time $t$ and $X(t)=1$ if a patient was diagnosed with disease by time $t$. This approach does not account for interval censoring, a major limitation in the cases considered in this study.
@@ 2.2 Piecewise-constant model accounting for interval-censoring
where $c_{k}$ are the times at which the hazard for transition $h\rightarrow l$ may change. This model is implemented in the msm package by Jackson, (2011) and can account for interval-censoring and the probability of developing disease between last disease scan and death or lost to follow-up. The values $c_{k}$ of the piecewise-constant intervals may differ for the different transitions.
The hazards for transitions to the death state can also be assumed proportional, allowing the effect of disease on survival to be modeled by constraining the transition hazards $\lambda_{12}$ and $\lambda_{02}$ to be equal and assuming the same piecewise-constant intervals. This can be done using the qconstrained option and by defining a binary time-dependent covariate $X(t)$, which equals 0 if the patient has not yet been diagnosed with the disease at time $t$, and 1 if the diagnosis has occurred by that time. The transition hazard from state 1 (disease) to state 2 (death) is then equal to $\lambda_{12}(t)=\lambda_{02}\exp\{\beta X(t)\}$.
@@ 2.4 M-spline model accounting for interval-censoring
The M-spline model offers greater flexibility than parametric models, but it also requires careful handling of the smoothing parameters in the penalized likelihood to ensure stable estimation. To this purpose, the idm function provides the option to automatically select the smoothing parameters using approximate leave-one-out cross-validation (CV=1) (Touraine et al., , 2017) However, cross-validation can be computationally intensive, and manual tuning of the smoothing parameters may be preferred, which can be challenging in complex datasets and is not always suitable for simulation studies.
@@ 3 Discrimination for time-dependent marker and survival data
To extend the concepts of sensitivity and specificity to settings with censored data and longitudinal markers, several definitions of cases and controls have been proposed. In this article, two of these definitions are considered: 1) incident cases and dynamic controls, and 2) cumulative cases and dynamic controls (Heagerty et al., , 2000; Heagerty and Zheng, , 2005; Zheng and Heagerty, , 2007). The two approaches target distinct aspects of marker performance: incident/dynamic metrics evaluate how well the marker discriminates individuals who will experience death at a specific time point among those still at risk, whereas cumulative/dynamic metrics assess how well the marker discriminates individuals who will experience death within a given time horizon.
@@ 3.1 Incident cases and dynamic controls
$p(t)=$ $\Pr\left(X_{i}(t-)=1\mid T^{D}_{i}=t\right)=\frac{P_{01}(t-)\lambda_{12}(t)}{P_{00}(t-)\lambda_{02}(t)+P_{01}(t-)\lambda_{12}(t)},$ (11)
@@
$\hat{\pi}_{1}(t)=\frac{\hat{P}_{01}(t)}{\hat{P}_{00}(t)+\hat{P}_{01}(t)}\quad\text{ and }\quad\hat{p}(t)=\frac{\hat{P}_{01}(t-)\frac{\hat{\lambda}_{12}(t)}{\hat{\lambda}_{02}(t)}}{\hat{P}_{00}(t-)+\hat{P}_{01}(t-)\frac{\hat{\lambda}_{12}(t)}{\hat{\lambda}_{02}(t)}},$
@@ 3.2 Cumulative cases and dynamic controls
where $T^{D}$ is time of death, $X(s)$ is marker measurement at time $s$. Cases are individuals who die within a time window (tst-s) from $s$ and controls are individuals who survive the time window. The cumulative/dynamic AUC is then defined by
@@
$\hat{\pi}_{1}(s,t)=\frac{\hat{P}_{01}(0,s)\hat{P}_{11}(s,t)}{\hat{P}_{00}(0,t)+\hat{P}_{01}(0,t)}\quad\text{ and }\quad\hat{p}(s,t)=\frac{\hat{P}_{01}(0,s)\hat{P}_{12}(s,t)}{\hat{P}_{00}(0,s)\hat{P}_{02}(s,t)+\hat{P}_{01}(0,s)\hat{P}_{12}(s,t)},$
@@ 4 Simulation study
To study the discrimination performance of an interval-censored binary disease marker on survival a simulation study was conducted. Incident/dynamic and cumulative/dynamic AUC were computed for the different estimation procedures presented in Section 2: the Cox model with time-dependent disease marker, which ignores interval-censoring, and the illness-death model with piecewise-constant, Weibull or M-spline transition hazards. For the piecewise-constant model, the four change points $c_{k}$ at which the hazard may change were considered at 6, 30, 60, and 90 months. For the M-spline model, the default of 7 knots per transition was used. Automatic selection of the smoothing parameter via approximate leave-one-out cross-validation was not implemented, as this approach may be numerically unstable in simulation studies; similarly, manual tuning was avoided to ensure consistency and comparability across simulated datasets. Multiple data scenarios were simulated (Section 4.1) and results from the different methods were compared (Section 4.2). R source code is available at http://github.com/mspreafico/auc-idmIC.
@@ 4.1 Data generation and methods
Data were generated from Weibull transition hazards with a common shape parameter $k$ and different rate parameters $\alpha_{01}$, $\alpha_{02}$, and $\alpha_{12}$, using the following procedure. Let $y$ denote the total follow-up length, with visits occurring at every $\tau$ time units, so that the time-dependent binary covariate $X(t)$ for illness can be observed at visits $t=0,\tau,2\tau,3\tau,\dots,y$. For each individual:
  1. 1.
    Draw $U_{1}\sim\text{Uniform}(0,1)$ and compute the Weibull-distributed first true exit time from state $h=0$ (disease-free) using the inversion method:
    $T_{1}=\left(\frac{-\ln U_{1}}{\alpha_{01}+\alpha_{02}}\right)^{1/k}.$
  2. 2.
    Decide which transition $h\rightarrow l$ occurred at $T_{1}$ by randomly sampling the first exit state $l\in\{1,2\}$ with probabilities
    $P(l=1)=\frac{\alpha_{01}}{\alpha_{01}+\alpha_{02}},\quad P(l=2)=\frac{\alpha_{02}}{\alpha_{01}+\alpha_{02}}.$
    If $l=2$, then the individual died at time $T^{D}=T_{1}$ without illness.
    If $l=1$, then the individual enters the disease state and $T_{1}$ is true illness time.
  3. 3.
    If the individual entered the disease state, draw $U_{2}\sim\text{Uniform}(0,1)$ and compute the true death time using the inversion method conditional on having already waited to $T_{1}$:
    $T^{D}=\left(T_{1}^{k}-\frac{\ln U_{2}}{\alpha_{12}}\right)^{1/k}.$
  4. 4.
    Censor the true death time according to the desired censoring scheme:
    • Administrative censoring: the survival time is $T^{*}=\min(T^{D},y)$ with death indicator $\delta=\mathbf{1}\{T^{D}\leq y\}$;
    • Uniform random censoring: draw the censoring time $C\sim\text{Uniform}(a,y)$ and determine the survival time as $T^{*}=\min(T^{D},C)$ with death indicator $\delta=\mathbf{1}\{T^{D}\leq C\}$.
  5. 5.
    Let $v^{*}=\max\{t\leq T^{*}:t\in\{0,\tau,2\tau,\dots,y\}\}$ denote the last observed visit before death or censoring. If the individual died without illness or was censored at $T^{*}<T_{1}$, then $X(t)=0$ for all $t=0,\tau,2\tau,\dots,v^{*}$. For individuals that entered the disease state before dying or being censored ($T^{*}\leq T^{1}$), map the true illness time $T_{1}$ to the first scheduled visit at/after illness and determine:
[…]
    where $v=\min\{t\geq T_{1}:t\in\{0,\tau,2\tau,\dots,v^{*}\}\}$ denotes the first visit after $T_{1}$, i.e., the observed illness time. In both cases $X(T^{*})=X(v^{*})$.
  6. 6.
    Return the time-to-death outcome $(T^{*},\delta)$, and the time-dependent covariate $\left\{X(t):t\in\{0,\tau,\dots,v^{*},T^{*}\}\right\}$ for observed disease.
@@
Table 1 summarizes the 18 scenarios (A–R) generated using shape parameter $k=0.5$ and rate parameters $\alpha_{01}=\alpha_{02}=0.05$ and $\alpha_{12}=0.56$. These Weibull parameters were kept fixed throughout the simulated scenarios and were based on the data discussed in Section 5, assuming a proportional hazards between diseased and non diseased with HR equal to 11.2. The number of individuals per data set, $N$, was either 400, 1000 or 2000. The survival time was censored according to two different censoring schemes: either it was censored administratively at 10 years ($y=120$ months) follow-up or censoring times were sampled from a uniform distribution between 5 and 10 years ($a=60$ and $y=120$ months). The disease state was observed only at pre-specified follow-up visits. The scenarios cover three different follow-up schemes in which the disease state was observed every 3, 6, or 12 months ($\tau=3,6,12$).
@@
The true incident/dynamic and cumulative/dynamic AUC values over time were determined using Equations (9) and (15), where the terms $\pi_{1}(t)$ and $p(t)$ were computed from the transition probabilities and hazards derived from the Weibull illness-death model with $k=0.5$, $\alpha_{01}=\alpha_{02}=0.05$, and $\alpha_{12}=0.56$, as detailed in Supplementary Material B. Each scenario was simulated $n_{sim}=1000$ times. For each model and scenario, estimated incident/dynamic and cumulative/dynamic AUCs over the years were presented graphically. Estimated bias, empirical standard error, and root mean squared error (Morris et al., , 2019) of both AUCs at specific time-points were considered as performance measures.
@@
Table 1: Simulated scenarios. NN is the total number of patients. Censoring is the type of death censoring: Unif(5, 10) means censoring was uniformly sampled between 5 and 10 years and 10 means that administrative censoring occurred at 10 years. Follow-up is time between disease observations in months.
Scenario NN Censoring Follow-up
@@ 4.2 Results
Hazard ratios for disease (yes vs no) were estimated for the piecewise-constant and the Cox model. Average values over $n_{sim}=1000$ repetitions were summarized in Supplementary Table C.1. For the Weibull and M-spline model no effect could be estimated, since the idm function does not allow transition hazards to be proportional. The coefficients from the Cox model were consistently more biased than the ones from the piecewise-constant model. The Cox model underestimated the true coefficient and the bias increased for larger follow-up intervals. Empirical SE decreased as sample size increased and was comparable between the Cox and piecewise-constant models, as well as across scenarios with equal sample sizes. These results are in line with Leffondré et al., (2013) who showed that the effect estimates of the Cox model were biased if the covariate affected both the risk of disease and death.
The M-spline model did not converge for many simulated data sets, with the frequency of invalid estimations (Supplementary Table C.2) increasing as sample size decreased. These difficulties are partly attributable to the decision not to implement automatic smoothing parameter selection due to potential numerical instability in simulations and to avoid manual tuning to ensure comparability across datasets. These invalid estimates prevented the estimation of the incident/dynamic and cumulative/dynamic AUCs. Therefore, the results for the M-spline model shown in Sections 4.2.2 and 4.2.3 are only based on the obtained valid estimates.
@@ 5.1 Soft tissue sarcoma data
Figure 4: Soft tissue sarcoma illness-death model (NN = 2232). State 0 (disease-free): surgery with curative intent; state 1 (disease): development of distant metastases (DM); state 2: death. The numbers indicate the number of patients moving from one state to the other.
@@ 5.2 Results
The HRs estimated by the Cox model and the piecewise-constant model were equal to 11.71 (95% CI = [10.31; 13.29]) and 11.28 (95% CI = [9.82; 12.96]) respectively. For the M-spline model, $m_{hl}=5$ knots were used for each transition, and the smoothing parameters of the penalized likelihood were selected using the approximate leave-one-out cross-validation procedure implemented in the SmoothHazard package, by setting idm(..., CV = 1) (Touraine et al., , 2017). However, the M-spline model converged only after rescaling the time variable, whereas the parametric Weibull model converged without any rescaling. Figure 5 displays on the left panel the non-parametric cumulative baseline hazards and on the right panel a graphical check of their fit to a Weibull distribution. For this figure the time of DM was assumed to be equal to the time that DM was detected during screening. If the data came from a Weibull distribution, the plot of the log-cumulative hazard against the log of time should be approximately linear (right panel), which is not the case in particular for the transition from surgery to DM (black line). The Weibull model therefore may not be appropriate for this data.
@@
Figure 6 shows the AUCs over time for the different models (left panel: incident/dynamic; right panel: cumulative/dynamic), with time-specific values at years $t=1,2,3,4,5$ shown in Table 4. The incident/dynamic AUCs of the Weibull (orange) and M-spline (yellow) models are initially much higher than those of the other models and decline over time, suggesting better prognostic ability in the short term than in the long term. The incident/dynamic AUC of the piecewise-constant model (magenta) is the lowest among all methods up to 5 years, after which the M-spline performs worse. The cumulative/dynamic AUC of the Cox model (blue) is the largest from $t=1.5$ years after surgery onwards, while the Weibull model performs worst, except at the beginning of follow-up. The M-spline shows good performance in the short term but deteriorates over time.
@@
Cox ROC (green): estimate based on Cox model (4) through risksetAUC function (only for incident/dynamic AUC). Cox Prob (blue): estimate based on Cox model through mstate transition probabilities. PW-const (magenta): estimate based on piecewise-constant model (5) through msm. Weibull (orange): estimate based on Weibull model (6) through SmoothHazard. M-spline (yellow): estimate based on M-spline model (7) with $m_{hl}=5$ knots per transition through SmoothHazard.
@@
Time-specific AUC Year $t$ Cox ROC Cox Prob PW-const Weibull M-spline
$\widehat{\text{AUC}}^{\text{I/D}}(t)$ 1 0.74 0.75 0.71 0.81 0.83
2 0.76 0.76 0.73 0.78 0.81
3 0.76 0.76 0.73 0.76 0.78
4 0.75 0.75 0.71 0.74 0.73
5 0.74 0.74 0.70 0.72 0.69
$\widehat{\text{AUC}}^{\text{C/D}}(t,t+5)$ 1 0.64 0.62 0.62 0.65
2 0.69 0.66 0.63 0.67
3 0.70 0.68 0.64 0.65
4 0.68 0.66 0.64 0.63
5 0.67 0.66 0.64 0.61
Note: Cox ROC, estimate based on Cox model (4) through risksetAUC function (only for incident/dynamic AUC); Cox Prob, estimate based on Cox model through mstate transition probabilities; PW-const, estimate based on piecewise-constant model (5) through msm; Weibull, estimate based on Weibull model (6) through SmoothHazard. M-spline, estimate based on M-spline model (7) with $m_{hl}=5$ knots per transition through SmoothHazard.